OpenStudy (anonymous):

Solve each equation. -4(x+9) = -12x-4

OpenStudy (anonymous):

-4x-36=-12x-4 8x=32 x=4

OpenStudy (luigi0210):

Factor out the -4(x+9): -4x-36=12x-4 --------------------------------- Move everything to one side: -16x=32 --------------------------------- Divide: x=-2

OpenStudy (anonymous):

Wait, did I look at that problem entirely wrong?

OpenStudy (luigi0210):

Oops, did not see that negative

OpenStudy (luigi0210):

and no you got it right smokey :P I just mis-read the problem

OpenStudy (anonymous):

Ah okay. I was beginning to lose faith in all math after that... And I already got through Calc C...

OpenStudy (luigi0210):

Calculus C?

OpenStudy (anonymous):

Yeah, that would have all been a waste if I couldn't even do simple algebra...

OpenStudy (luigi0210):

You're a smart girl I'm sure :P